An electron initially has a speed 19.4 km/s along the x-direction and enters an electric field of strength 24 mV/m that points in the y-direction and has a width of 18.5 cm. How far is the deflection?
1
Expert's answer
2019-09-16T09:46:17-0400
"d=vt"

"t=\frac{d}{v}=\frac{0.185}{19400}=9.536\cdot 10^{-6}\ s"


The electrostatic force:


"F=qE"

The acceleration:


"a=\frac{F}{m}=\frac{qE}{m}"

The deflection is


"h=0.5at^2=\frac{qE}{2m}t^2"

"h=\frac{(1.6\cdot 10^{-19})(0.024)}{2(9.11\cdot 10^{-31})}(9.536\cdot 10^{-6})^2=0.192\ m"